OpenStudy (anonymous):

use the substitution method to find the value of y in the given system of equations. express non-integer as fractions in lowest terms. x=3y 2x -5y= 1/8 y=_________

OpenStudy (anonymous):

2(3y)-5y=1/8 6y-5y=1/8 y=1/8

OpenStudy (anonymous):

x=3(1/8) x=3/8